# Semantic Interaction Information mediates compositional generalization in latent space

### 28 Mar 2026 v1

@@ 3.4 Experiment 3: Latent space navigation enables offline optimization
Figure 7: Conditional generative modeling enables optimization in compositional spaces. (a) Schematic illustrating the mapping of preferred observations ($\mathbf{\Omega}$) to their respective likelihoods and the cumulative landscape (accumulated over $i$ in subsection 3.4). (b) An example traversal, from the lowest to the highest point on the landscape, changes observations to best match the agent’s preference. (c) Controller learning curves (mean $\pm$ SEM, $n=20$ initializations of $\mathbf{\Omega}$). Performance is evaluated as $\mathcal{R}(\text{argmax}_{\mathbf{r}}\pi(\mathbf{r}),\mathbf{\Omega},\ell,Z)/\max_{\mathbf{r}}\mathcal{R}(\mathbf{r},\mathbf{\Omega},\ell,Z)$ (using the true $\ell$ and $Z$).
The Controllers are trained through a standard Advantage Actor-Critic (A2C) algorithm. We compare training Controllers across three conditions that vary estimates of $\hat{Z}$ and $\hat{\ell}$:
@@
Training offline with the generative model achieves the same level of performance as training online or with access to the joint likelihood itself (Figure 7c). Figure 8 displays several example trajectories of the policy’s maximum likelihood realization, $\text{argmax}_{\mathbf{r}}\pi(\mathbf{r})$, for the Offline agent w/ Generator. These traversals reveal the utility of compositional embeddings for offline optimization.
Figure 8: Example offline learning trajectories w/ Generator. The evolution of the deterministic policy, $\text{argmax}_{\mathbf{r}}\pi(\mathbf{r})$, is plotted throughout offline training from initialization (red circles) to the end of training (green stars). These trajectories are overlaid on the preference landscapes to demonstrate navigation in the internally generated Cognitive Gridworld.
@@ 4 Discussion
In this work, we operationalized continual meta-learning as the conjunction of variational and sequential inference. This allowed us to establish the Cognitive Gridworld as a testbed for evaluating generalization from goal-directed experience. Building off previous work, we demonstrated that Representation Classification Chains enable an agent to learn compositional embeddings from natural feedback and passive experience. These embeddings can then be utilized to optimize novel objectives without further environmental interaction. It is notable that while Representation Classification Chains were trained on individual training states, compositional embeddings enabled the Offline Controller to optimize a policy in a multi-dimensional (joint) testing state space. This builds upon [2]’s work, recasting model-free RL as a conditional generative modeling problem, which enabled generalization to a novel combination of inputs (e.g. constraints, reward functions, etc). Conversely, by recasting latent space traversals as a model-free RL problem, RCC’s generalize to novel combinations of outputs (e.g. decision variables, control variables etc). This may be particularly helpful for optimizing non-differentiable objectives. In summary, RCCs serve as promising evaluators that learn passively during goal-directed experience. Their ability to evaluate in imagined latent spaces allows for zero-shot generalization to novel combinations of relevant causal factors. Future work should attempt to integrate RCCs with the VASE algorithm to identify which embeddings are relevant to a task, rather than assuming this is known.
@@ 5.1 Experiment 1:
The goal belief $B^{\text{Net}}_{tgr}$ is extracted from the accumulated output $\mathcal{M}_{tcr}=\sum_{t^{\prime}=1}^{t}M_{t^{\prime}cr}$ (where $M_{t^{\prime}cr}$ is single timestep network pre-activations) via a softmax over possible realizations $B^{\text{Net}}_{tgr}=\frac{e^{\mathcal{M}_{tgr}}}{\sum_{r^{\prime}}e^{\mathcal{M}_{tgr^{\prime}}}}.$ Although observations are conditionally independent given the joint latent state $(r_{g},r_{c})$, the optimal maximum a posteriori (MAP) estimate of $r_{g}$ requires marginalizing the joint likelihood over $r_{c}$ before maximizing:
@@ 5.2 Experiment 2:
We sample a latent configuration $\hat{\mathbf{r}}$ from the Classifier belief at $t=T$ (without a gradient), and train the Generator to maximize the likelihood of observations. Since the true generative distribution is unknown, we construct an empirical target distribution $\tilde{P}(\mathbf{o})$, where the likelihood of each independent observable is given by its trajectory average $\tilde{P}(o^{i}=1)=\langle o^{i}\rangle_{t}$. We use this to compute a symmetric $\mathcal{D}_{\mathrm{KL}}$ (hence a self-supervised loss):
@@ Appendix A SUPPLEMENTARY
A.1 Defining compositional generalization in latent space
In our setting, Compositionality refers to the interaction between variables, where interaction is operationalized as a cosine similarity between embeddings associated with each variable. Generalization refers to indirectly learning (from partial feedback) variable embeddings in a manner which will generalize to a novel combination of variables. Partial feedback is operationalized by Goals, which refer to a subset of variables that receive a teaching signal during training. Latent Space refers to possible values of the random variables, with a specific realization reflecting a specific state of the world. In summary, compositional generalization in latent space refers to indirectly learning how a set of observations were generated by a novel combination of relevant state space dimensions.
A.2 Generative process details
Figure 9: A flexible process for embedding Gridworld structure into latent space. The embeddings of active states are compressed into interactions which are then expanded to a set of vectors that map to a discrete probability distribution over possible realizations of the world.
@@ A.2 Generative process details
Generalization emerges when novel set $\{\mathbf{k}^{i}_{c},\mathbf{q}_{c}^{i},\dots,\mathbf{k}^{i}_{{C}},\mathbf{q}_{{C}}^{i}\}$ reduces to seen $\mathbf{z}^{i}$. The low rank structure of $\mathbf{z}^{i}\in\mathbb{R}^{C\times(C-1)}$ is embedded in joint likelihood, $\ell_{\mathbf{z}^{i}}\in[0,1]^{{R}^{C}}$. Below we step through the default expansion from $\mathbf{z}^{i}\to\ell_{\mathbf{z}^{i}}$ in 2 steps.
@@
First, we expand each scalar $z^{i}_{cc^{\prime}}$ to vector $\mathbf{v}^{i}_{cc^{\prime}}$ through:
@@
Step 2. Expanding $\mathbf{v}_{cc^{\prime}}^{i}(\mathbf{v}_{cc^{\prime}}^{i})^{\top}\to\ell_{\mathbf{z}^{i}}\in[0,1]^{R^{C}}$:
Next, we expand outer products $\mathbf{v}_{cc^{\prime}}^{i}(\mathbf{v}_{cc^{\prime}}^{i})^{\top}$ to tensor $\ell_{\mathbf{z}^{i}}\in[0,1]^{R^{C}}$ through:
@@
$\ell_{\mathbf{z}^{i}}=\sigma\bigl(\{\mathbf{v}^{i}_{12}(\mathbf{v}^{i}_{21})^{\top}\}_{12}+\{\mathbf{v}^{i}_{13}(\mathbf{v}^{i}_{31})^{\top}\}_{13}+\{\mathbf{v}^{i}_{23}(\mathbf{v}^{i}_{32})^{\top}\}_{23}\bigr)$
@@
  • Useful Baseline: Ensuring the potential for asymmetry in $\mathbf{v}^{i}_{cc^{\prime}}$ ensures that independent inference can perform above chance.
@@
A.3 SII Relation to Interaction Information
Interaction Information, $I(r_{g};r_{c};o)$, quantifies how the mutual information between two random variables, $r_{g}$ and $r_{c}$, changes when conditioned on a third, $o$. Assuming the latent variables $r_{g}$ and $r_{c}$ are a priori independent ($I(r_{g};r_{c})=0$) (as they are in the setting we consider), the definition simplifies to (the negative of) conditional mutual information, or equivalently, the (negative) expected ($\mathcal{D}_{\mathrm{KL}}$) between the conditional joint distribution and the product of its conditional marginals:
$I(r_{g};r_{c};o)=-I(r_{g};r_{c}\mid o)=-\mathbb{E}[\mathcal{D}_{\mathrm{KL}}(p(r_{g},r_{c}\mid o)\,||\,p(r_{g}\mid o)\otimes p(r_{c}\mid o))]$
Note: The marginals of $p(r_{g},r_{c}\mid o)$ are not necessarily the posteriors under independent inference.
Let us now consider two types of Bayesian observers processing a sequence of observations $o_{1:T}$:
@@ A.3 SII Relation to Interaction Information
    Optimal Observer: Updates its joint belief $B_{t}(r_{g},r_{c})$ correctly using $p(o|r_{g},r_{c})$.
@@
    $\tilde{p}(o|r_{g},r_{c})\propto p(o\mid r_{g})\otimes p(o\mid r_{c})$
@@
$\sum_{r_{c}}p(o_{1}|r_{g},r_{c})p(r_{g},r_{c})=\sum_{r_{c}}p(o_{1}\mid r_{g})\otimes p(o_{1}\mid r_{c})p(r_{g},r_{c})$
Therefore, for the Interaction Information produced by a single observation ($o_{1}$), it is valid to replace the conditional distributions with our Optimal and Naive observers:
$I(r_{g};r_{c};o_{1})=-\mathbb{E}[\mathcal{D}_{\mathrm{KL}}(B_{0}(r_{g},r_{c}|o_{1})\,||\,\tilde{B}_{0}(r_{g}|o_{1})\otimes\tilde{B}_{0}(r_{c}|o_{1}))]$
@@
$SII_{T}=\mathbb{E}[\mathcal{D}_{\mathrm{KL}}(B_{T}(r_{g}|o_{1:T})\,||\,\tilde{B}_{T}(r_{g}|o_{1:T}))]$
Equivalently, SII measures the residual Interaction Information in $o_{1:T}$ after conditioning on $r_{g}$:
$SII_{T}\;\propto\;\mathbb{E}\!\Big[\mathcal{D}_{\mathrm{KL}}\!\big(p(o_{1:T}|r_{g})\;\|\;\prod_{t}p(o_{t}|r_{g})\big)\Big],$
@@
Summary: $SII$ is not $I(r_{g};r_{c};o)$, rather it is $I(o_{1},\dots,o_{T},r_{g})$, a measure of the accumulated, reward-relevant information which would be lost by neglecting interaction information. SII is the cost of a myopic focus on the goal-pertinent state and distinguishes a genuine world model from strategies that simply exploit spurious correlations.
@@
A.4 Learning, Dynamics and Computation
@@
Figure 11: Early $C=2$ learning dynamics reveal different strategies. (a) Average final step accuracy improves throughout training for both Fully Trained and Echo State networks during early training. (b) A negative correlation between accuracy and Semantic Interaction Information (SII) emerges early on, and persists for the Echo State network. (c) The predictability of accuracy from SII closely tracks the Participation Ratio (PR) of the recurrent dynamics minus the Participation Ratio of the read-in activity. Over time, the Fully Trained network expands its recurrent dimensionality, while the Echo State network develops a more expressive read-in.
Future theoretical work is still needed to specify the relationship between learning, dynamics and computation. For instance, we observed that while average performance improves throughout training for both Fully Trained and Echo State networks (Figure 11a), a negative correlation between accuracy and Semantic Interaction Information emerges early on (Figure 11b). This relationship continues to strengthen for an Echo State network, yet decays when recurrent weights are plastic. This indicates that while average performance improves, the network solution may be moving further from optimality. A natural analogy is the exploration-exploitation tradeoff, where finetuning a reliable strategy could lead the agent away from discovering a superior strategy. One possibility is that the reliable strategy, when learning to compute, is externally-driven dynamics that amplify confidence in the typical case. Consistent with this possibility, we found that the predictability of accuracy from SII has a strong correspondence with the relative Participation Ratio of the read-in activity to the recurrent dynamics (Figure 11c). While a Fully Trained network expands the dimensionality of its recurrent dynamics, the Echo State network is developing a more expressive encoder. Although this strategy initially improves accuracy, it also creates proactive interference, blocking the ability to learn subtle differences in a trajectory that correspond to large jumps in latent space.
@@
Misalignment between the RNN’s state space and the true latent space could be explained mechanistically as entanglement. Entanglement in an RNN’s representation may emerge from the mixing of information by temporal correlations between neurons [59]. Therefore, we hypothesized that training only a read-in and read-out on an evidence accumulation task, which demands temporal correlations, would not only fail to overcome this entanglement, but would exacerbate it. Indeed, Principal Component Analysis (PCA) on the marginal beliefs for network and Bayesian agents was consistent with this prediction. When performing PCA on marginal beliefs after a single observation, all representations appeared factorized (Figure 12). Yet PCA on the marginal beliefs at the end of inference appeared entangled for the Echo State and Naive Bayes (Figure 13) who lose the ability to differentiate $r_{c}$ from $r_{c^{\prime}}$. Additionally, the Joint and Fully Trained representation allocates low entropy beliefs (i.e. high confidence) to all $(r_{c},r_{c^{\prime}})$ pairs. In contrast, the Echo State network and Naive Bayes represent confidence in an orthogonal direction to $r_{c}$ and $r_{c^{\prime}}$.
Figure 12: Belief representations are initially factorized. Top 2 Principal Components of the marginal beliefs at the first time step ($t=1$) for Fully Trained Networks, Echo State networks, Joint Bayes and Naive Bayes. Beliefs are colored by the realization sum $r_{c}+r_{c^{\prime}}$ (top), difference $r_{c}-r_{c^{\prime}}$ (middle), and average belief entropy $\sum_{c}\sum_{r}B_{1cr}\ln B_{1cr}$ (bottom).
@@
Figure 13: Failing to capture interaction information causes entanglement over time. Same as Figure 12, for the final observation ($t=T$). The Fully Trained network successfully maintains factorized representations akin to Joint Bayes. In contrast, the Echo State and Naive Bayes representations of $r_{c}$ and $r_{c^{\prime}}$ become entangled, as evidenced by the inability to linearly separate them from the top 2 components.
@@
The following section provides a potential source of this link between the entanglement in mechanistic and analytical solution. Let $\mathbf{e}_{r}\in\{0,1\}^{R}$ represent a one-hot indicator vector which inquires about the probability of realization $r$. In $P_{z^{i}}(o^{i}\mid r)$, the mapping from scalar $z^{i}$ to a vector of $R$ updates is mediated by expansion $f(z^{i})$. For $C=1$, inquiring about a specific realization $r=r^{\prime}$ projects the indicator vector onto the likelihood via:

### 20 Apr 2026 v3

@@ 3.4 Experiment 3: Latent space navigation enables offline optimization
Figure 7: Conditional generative modeling enables optimization in compositional spaces. (a) Schematic illustrating the mapping of preferred observations ($\mathbf{\Omega}$) to their respective likelihoods and the cumulative landscape (accumulated over $i$ in subsection 3.4). (b) An example traversal, from the lowest to the highest point on the landscape, changes observations to best match the agent’s preference. (c) Controller learning curves (mean $\pm$ SEM, $n=20$ initializations of $\mathbf{\Omega}$). Performance is evaluated as $\mathcal{R}(\text{argmax}_{\mathbf{r}}\pi(\mathbf{r}),\mathbf{\Omega},\ell,Z)/\max_{\mathbf{r}}\mathcal{R}(\mathbf{r},\mathbf{\Omega},\ell,Z)$ using the ground truth $\ell$ and $Z$.
The RL agents are trained through a standard Advantage Actor-Critic (A2C) algorithm. We compare Controller training across three conditions that vary in how they estimate $\hat{Z}$ and $\hat{\ell}$:
@@
Training fully offline with the generative model achieves the similar performance and efficiency as training offline with direct access to the joint likelihood and training online with direct interactions with the environment (Figure 7c). It is worth noting that the rapid learning seen in Figure 7c required a large batch of 20,000 trajectories per episode, each followed by an interaction with the environment in the Online condition. In the Offline conditions, training did not require any interactions with the environment. Figure 8 displays several examples of the Offline w/ Generator agent’s deterministic policy ($\text{argmax}_{\mathbf{r}}\pi(\mathbf{r})$) throughout several learning trajectories. These traversals reveal the utility of compositional embeddings for offline optimization. More representative landscapes and trajectories can be found in supplementary Figure 15.
Figure 8: Example offline learning trajectories w/ Generator. The evolution of the deterministic policy, $\text{argmax}_{\mathbf{r}}\pi(\mathbf{r})$, is plotted throughout offline training from initialization (red circles) to the end of training (green stars). Trajectories are overlaid on the preference landscapes to demonstrate navigation through an internal Cognitive Gridworld.
@@ 4 Discussion
Building off previous work, we demonstrated that Representation Classification Chains enable an agent to learn variable embeddings from natural feedback and passive experience. This novel architecture considers continual meta-learning as the conjunction of variational and sequential inference, continually learning embeddings on a separate timescale from the dynamics of inference. The learned embeddings can then be utilized to optimize novel objectives without further environmental interaction. It is notable that while Representation Classification Chains were trained on individual training variables, compositional embeddings enabled the Offline Controller to optimize a policy in a multi-dimensional joint action space of testing variables. This extends the work of [2], which cast model-free RL as a conditional generative modeling problem to enable compositional generalization in the input space. Conversely, by recasting latent space traversals as a model-free RL problem, RCCs generalize to novel combinations of outputs (e.g. decision variables, control variables etc). This may be particularly helpful for optimizing non-differentiable objectives. Future work should attempt to integrate RCCs with the VASE algorithm to identify which embeddings are relevant to a task, rather than assuming this is known.
@@ 5.1 Experiment 1:
The goal belief $B^{\text{Net}}_{tgr}$ is extracted from the accumulated output $\mathcal{M}_{tcr}=\sum_{t^{\prime}=1}^{t}M_{t^{\prime}cr}$ (where $M_{t^{\prime}cr}$ is single timestep network pre-activations) via a softmax over possible realizations $B^{\text{Net}}_{tgr}=\frac{e^{\mathcal{M}_{tgr}}}{\sum_{r^{\prime}}e^{\mathcal{M}_{tgr^{\prime}}}}.$ Although observations are conditionally independent given both latent variable realizations $(r_{g},r_{c})$, the optimal maximum a posteriori (MAP) estimate of $r_{g}$ requires marginalizing the joint likelihood over $r_{c}$ before maximizing:
@@ 5.2 Experiment 2:
We sample a latent configuration $\hat{\mathbf{r}}$ from the Classifier belief at $t=T$ (without a gradient), and train the Generator to maximize the likelihood of observations. Since the true generative distribution is unknown, we construct an empirical target distribution $\tilde{P}(\mathbf{o})$, where the likelihood of each independent observable is given by its trajectory average $\tilde{p}_{Z}(o^{i}=1)=\langle o^{i}\rangle_{t}$. We use this to compute a symmetric $\mathcal{D}_{\mathrm{KL}}$ (hence a self-supervised loss):
@@ Appendix A SUPPLEMENTARY
A.1 Generative process details
Figure 9: A flexible process for embedding Gridworld structure into latent space. The embeddings of relevant variables are compressed into interactions which are then expanded to a discrete probability distribution over possible realizations of the world. The full process consists of first (i) compressing embedding vectors to their scalar interactions. Then (ii) expanding pairwise interactions to pairs of vectors, expanding the vector pairs to matrices through an outer product, and (for $C\geq 3$) expanding the matrices into a tensor by broadcasting each matrix over all other dimensions. Finally, a valid (iii) likelihood is ensured by a sigmoid transformation for normalization.
@@ A.1 Generative process details
Extrapolation to an out of distribution set $\{\mathbf{k}^{i}_{c},\mathbf{q}_{c}^{i},\dots,\mathbf{k}^{i}_{{C}},\mathbf{q}_{{C}}^{i}\}$ may occur if the resultant $\mathbf{z}^{i}$ falls within the manifold of training $\mathbf{z}^{i}$. The low rank structure of $\mathbf{z}^{i}\in\mathbb{R}^{C\times(C-1)}$ is embedded in joint likelihood, $\ell_{\mathbf{z}^{i}}\in[0,1]^{{R}^{C}}$. Below we step through the default expansion from $\mathbf{z}^{i}\to\ell_{\mathbf{z}^{i}}$ in 2 steps.
@@
First, we expand each scalar $z^{i}_{cc^{\prime}}$ to vector $\mathbf{v}^{i}_{cc^{\prime}}$ through $\Theta$ and $\omega$:
@@
Step 2. Expanding $\mathbf{v}_{cc^{\prime}}^{i}(\mathbf{v}_{c^{\prime}c}^{i})^{\top}\to\ell_{\mathbf{z}^{i}}\in[0,1]^{R^{C}}$:
Next, we expand outer products $\mathbf{v}_{cc^{\prime}}^{i}(\mathbf{v}_{c^{\prime}c}^{i})^{\top}$ to tensor $\ell_{\mathbf{z}^{i}}\in[0,1]^{R^{C}}$ through:
@@
$\ell_{\mathbf{z}^{i}}=\sigma\bigl(\{\mathbf{v}^{i}_{12}(\mathbf{v}^{i}_{21})^{\top}\}_{12}+\{\mathbf{v}^{i}_{13}(\mathbf{v}^{i}_{31})^{\top}\}_{13}+\{\mathbf{v}^{i}_{23}(\mathbf{v}^{i}_{32})^{\top}\}_{23}\bigr).$
@@
  • Useful Baseline: Ensuring the potential for asymmetry in $\mathbf{v}^{i}_{cc^{\prime}}$ ensures that naive inference can perform above chance.
@@
A.2 SII Relation to Interaction Information
Interaction Information, $I(r_{g};r_{c};o^{i})$, quantifies how the mutual information between two random variables, $r_{g}$ and $r_{c}$, changes when conditioned on a third, $o^{i}$. Assuming the latent variables $r_{g}$ and $r_{c}$ are a priori independent ($I(r_{g};r_{c})=0$) (as they are in the setting we consider), the definition simplifies to (the negative of) conditional mutual information, or equivalently, the (negative) expected $\mathcal{D}_{\mathrm{KL}}$ between the conditional joint distribution and the product of its conditional marginals:
$I(r_{g};r_{c};o^{i})=-I(r_{g};r_{c}\mid o^{i})=-\mathbb{E}[\mathcal{D}_{\mathrm{KL}}(p_{Z^{i}}(r_{g},r_{c}\mid o^{i})\,||\,p_{Z^{i}}(r_{g}\mid o^{i})\otimes p_{Z^{i}}(r_{c}\mid o^{i}))]$
Note that the marginals of $p_{Z^{i}}(r_{g},r_{c}\mid o^{i})$ are not the posteriors of independent inference.
Let us now consider two types of Bayesian observers processing a sequence of observations $o^{i}_{1:T}$:
@@ A.2 SII Relation to Interaction Information
    Optimal Observer: Updates its joint belief $B_{t}(r_{g},r_{c})$ correctly using $p_{Z^{i}}(o^{i}|r_{g},r_{c})$.
@@
    $\tilde{p}_{Z^{i}}(o^{i}|r_{g},r_{c})\propto p_{Z^{i}}(o^{i}\mid r_{g})\otimes p_{Z^{i}}(o^{i}\mid r_{c})$
@@
$\sum_{r_{c}}p_{Z^{i}}(o^{i}_{1}|r_{g},r_{c})p(r_{g},r_{c})=\sum_{r_{c}}p_{Z^{i}}(o^{i}_{1}\mid r_{g})\otimes p_{Z^{i}}(o^{i}_{1}\mid r_{c})p(r_{g},r_{c})$
Therefore, for the Interaction Information produced by a single observation ($o^{i}_{1}$), it is valid to replace the conditional distributions with our Optimal and Naive observers:
$I(r_{g};r_{c};o^{i}_{1})=-\mathbb{E}[\mathcal{D}_{\mathrm{KL}}(B_{0}(r_{g},r_{c}|o^{i}_{1})\,||\,\tilde{B}_{0}(r_{g}|o^{i}_{1})\otimes\tilde{B}_{0}(r_{c}|o^{i}_{1}))]$
@@
$SII_{T}=\mathbb{E}[\mathcal{D}_{\mathrm{KL}}(B_{T}(r_{g}|o^{i}_{1:T})\,||\,\tilde{B}_{T}(r_{g}|o^{i}_{1:T}))]$
Equivalently, SII measures the residual Interaction Information in $\mathbf{o}_{1:T}$ after conditioning on $r_{g}$:
$SII_{T}\;\propto\;\mathbb{E}\!\Big[\mathcal{D}_{\mathrm{KL}}\!\big(p_{Z}(\mathbf{o}_{1:T}|r_{g})\;\|\;\prod_{t}\prod_{i}p_{Z^{i}}(o^{i}_{t}|r_{g})\big)\Big],$
@@
Summary: $SII$ is not $I(r_{g};r_{c};\mathbf{o}_{t})$, rather it is $I(\mathbf{o}_{1},\dots,\mathbf{o}_{T},r_{g})$, a measure of the accumulated, reward-relevant information which would be lost by neglecting interaction information. SII is the cost of a myopic focus on the goal-pertinent variable and distinguishes a genuine world model from strategies that simply exploit spurious correlations.
@@
A.3 Learning Dynamics
Future theoretical work is still needed to specify the relationship between learning, dynamics and computation. For instance, we observed that while average performance improves throughout training for both Fully Trained and Echo State networks (Figure 11a-b), a negative correlation between accuracy and Semantic Interaction Information emerges early on (Figure 11c). It is important to highlight this critical finding that remains poorly understood.
Since the network initially knows nothing, the correlation between SII and accuracy begins at zero. Yet rapidly, failure becomes predictable from SII because the agent is failing in trickier trajectories that require second guessing Naive Bayes. Naivety may be conceptualized as a short-term reward which must be actively resisted for nuance, akin to an informational Marshmallow test [60]. Importantly, this correlation dissolves for the Fully Trained network, but persists when the recurrent weights are frozen. Despite the Echo State performance improving on average, its solution is moving further from optimality. A natural analogy is the exploration-exploitation tradeoff, where finetuning an early strategy could lead the agent away from discovering a superior strategy.
We hypothesize that this suboptimal strategy is stimulus bound [104]. To test this hypothesis, we investigate the Participation Ratio (i.e. dimensionality) of the Read-in activity and recurrent activity (Figure 11d). While a Fully Trained network expands the dimensionality of its recurrent dynamics, the Echo State network develops a more expressive encoder. Although this strategy improves accuracy, it may prevent the discovery of subtle changes in a trajectory that correspond to large jumps in latent space. Ultimately, we find that externally-driven dynamics are associated with a negative correlation between SII and accuracy. Despite early learning, the Echo State network remains likely to fail in trajectories with high SII, leading to an upper bound characteristic of Naive Bayes.
@@ A.3 Learning Dynamics
Figure 11: Early correlation between accuracy and SII predicts eventual performance ($C=2$). (a) Throughout learning, the testing accuracy at the final step of inference saturates to the performance of either Joint or Naive Bayes. (b) Final step testing accuracy during early training. (c) Correlation between Semantic Interaction Information (SII) and accuracy at the final step of inference. A negative correlation emerges in the beginning of training and persists for the Echo State network. (d) Difference in Participation Ratio (PR) between the recurrent and read-in activity.
A.4 Dis-entanglement
@@
Misalignment between the Echo State network’s internal space and the true latent space could be explained mechanistically as entanglement. Entanglement in an RNN’s representation may emerge from the mixing of information by temporal correlations between neurons [59]. Therefore, we hypothesized that training only a read-in and read-out on an evidence accumulation task, which demands temporal correlations, would not only fail to overcome this entanglement, but would exacerbate it. Indeed, Principal Component Analysis (PCA) on the marginal beliefs for network and Bayesian agents was consistent with this prediction. When performing PCA on marginal beliefs after a single observation, all representations appeared factorized (Figure 12). Yet PCA on the marginal beliefs at the end of inference appeared entangled for the Echo State and Naive Bayes (Figure 13) which lose the ability to differentiate $r_{c}$ from $r_{c^{\prime}}$. Additionally, the Joint and Fully Trained representations allocate low entropy beliefs (i.e. high confidence) to all $(r_{c},r_{c^{\prime}})$ pairs. In contrast, the Echo State network and Naive Bayes represent confidence in an orthogonal direction to $r_{c}$ and $r_{c^{\prime}}$.
Figure 12: Belief representations are initially factorized. Top 2 Principal Components of the marginal beliefs after a single observation. Beliefs are colored by the realization sum ($r_{c}+r_{c^{\prime}}$), difference ($r_{c}-r_{c^{\prime}}$), and belief entropy ($\sum_{c}\sum_{r}B_{tcr}\ln B_{tcr}$). $R^{2}$ indicates the variance explained of the respective variables by the top 2 components.
@@
Figure 13: Failing to capture interaction information causes entanglement over time. Same as Figure 12, for the final observation. The Fully Trained network maintains factorized representations akin to Joint Bayes. The Echo State and Naive representations of $r_{c}$ and $r_{c^{\prime}}$ become entangled.
@@
For a mechanistic account of this entanglement, let $\mathbf{e}_{r}\in\{0,1\}^{R}$ represent a one-hot indicator vector which inquires about the probability of realization $r$. In $P_{z^{i}}(o^{i}\mid r)$, the mapping from scalar $z^{i}$ to a vector of $R$ updates is mediated by expansion $f(z^{i})$. For $C=1$, inquiring about a specific realization $r=r^{\prime}$ projects the indicator vector onto the likelihood via: